OpenStudy (anonymous):

consider the region bounded by the graphs of f(x)=x^1/2, y=0, and x=2. find the volume of the soild formed by rotating the region about the x-axis

OpenStudy (amistre64):

ok.... y = sqrt(x) good

OpenStudy (amistre64):

between [0,2]? right

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

we find an area of rotation by taking the radius and adding it up as we go thru the interval (S) 2pi(f(x)) dx or: 2pi (S) f(x) dx ... right?

OpenStudy (anonymous):

i have no idea....

OpenStudy (amistre64):

lol .... ok... I can work with that :)

OpenStudy (anonymous):

is the graph bounded by x \[\in\] [0, 2]?

OpenStudy (amistre64):

if we take one section of this graph, we have a radius = to the value of f(x) at the point x right?

OpenStudy (anonymous):

\[x \in [0, 2]\] ?

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

rdoshi, yes

OpenStudy (anonymous):

alrighty...

OpenStudy (amistre64):

if we take this radius and integrate 2pi(r) we will get the area of this "slice" lets use an example we are familiar with ok?

OpenStudy (amistre64):

lets say the radius = 5; the area of a circle would be equal to pi(r^2) right?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

I understand that, i'm just confused on how to do it with this problem...because I just AM DUMB

OpenStudy (amistre64):

if we take the Circumference of a circle and "integrate" it we get this: (S) 2pi(r) dr 2pi (S) r dr 2pi (r^2/2) = pi(r^2) which gives us the area of our circle.... 25pi

OpenStudy (anonymous):

\[\int\limits_{0}^{2}\pi \sqrt{x}\] That should do it?

OpenStudy (amistre64):

what we do by integrating along an interval is taking each "slice" finding its area and then adding up all the slices to get the whole volume

OpenStudy (anonymous):

sorry sqrt(x) is squared for area

OpenStudy (anonymous):

so its just integral (pi * x) from 0 to 2?

OpenStudy (amistre64):

our radius is equal to y = x^(1/2) does that make sense to you?

OpenStudy (anonymous):

you find the equation for the cross sectional area of the solenoid that is formed by rotating the graph which is A(x) = pi (sqrt x)^2 then you just find the integral of that area along x for the volume

OpenStudy (anonymous):

but to find the integral do you find the anti derviative?

OpenStudy (amistre64):

yes... an integral and an antiderivative are the same thing...

OpenStudy (amistre64):

beth: whats the antiderivative of x^(1/2)

OpenStudy (anonymous):

nd then i just plug in 2 and 0 and solve right

OpenStudy (amistre64):

actually, 0 will equal 0 so you just have to solve for x=2

OpenStudy (anonymous):

you dont take integral of sqrt(x)

OpenStudy (anonymous):

because to find area you have to square the radius

OpenStudy (anonymous):

sqrt(x) ^ 2 is just x... so you find the integral of pi(x) from 0 to 2

OpenStudy (anonymous):

but i am trying to find volume not area...or does that even matter?

OpenStudy (amistre64):

volume is just 2d area.... add up all the volumes of the slices to get the 3d area.... same thing

OpenStudy (anonymous):

alright here's how it works.... integrating area will give you volume. in order to find area A = pi radius^2... radius in this case is sqrt(x)

OpenStudy (amistre64):

2pi (S) x^(1/2) dx [0,2] 2pi(x^3/2)(2/3) when x = 2 is all you need to find

OpenStudy (amistre64):

(4/3)pi 2^(3/2) = ??

OpenStudy (anonymous):

pi (S) (x^1/2)^2 dx from 0 to 2

OpenStudy (anonymous):

What rdoshi said. Your should only need integrals for this question. by using the formula below it will find you the volume. when rotating around x: \[V = \pi \int\limits_{b}^{a} y^{2} dx\] [therefore this means f(x) is squared] when rotating around the y: \[V = \pi \int\limits_{b}^{a} x^{2} dy\] [therefore this means you make x the subject of the function and then square it]

OpenStudy (anonymous):

so the final answer should be 2pi?

OpenStudy (amistre64):

we already went thru what we need to do to get the "area"... so follow along :)

OpenStudy (amistre64):

4 pi 2^(3/2) ------------- 3

OpenStudy (anonymous):

you guys are telling me two different things...i dont know which one is right haaha

OpenStudy (anonymous):

its 2pi

OpenStudy (anonymous):

for area you just exclude pi from the formula while taking into account that if the function goes into a negative area you use absolute values (seeing you can't have negative an area).

OpenStudy (anonymous):

yea but i am trying to find volume, but wait that doesnt matter right?

OpenStudy (amistre64):

the area of each slice is the antiderivative of 2pi(x^(1/2))

OpenStudy (anonymous):

i thought the area of each slice is the antiderivative of pi((x^1/2)^2)

OpenStudy (amistre64):

the antiderivative of 2pi(r) = 2pi(r^2)/2

OpenStudy (anonymous):

2piR is circumfrence not area

OpenStudy (anonymous):

pi R^2 is area

OpenStudy (amistre64):

why would it be the antiderivative of pi(x^(1/2)^2)?? that makes no sense

OpenStudy (anonymous):

R in this case is y = sqrt(x)

OpenStudy (amistre64):

rdoshi... you antiderive the circumference of a circle to get the formula for the area of a circle.... its just that simple

OpenStudy (anonymous):

anti deravative of 2pi R is pi R^2

OpenStudy (amistre64):

yes... and the R value is y = x^(1/2) BEFORE you get the antiderivative...

OpenStudy (anonymous):

that doesn't change anything

OpenStudy (amistre64):

it does in my mind :)

OpenStudy (anonymous):

you can substitute before or after... method of substitution

OpenStudy (amistre64):

I perfer before... makes my life easier :)

OpenStudy (anonymous):

nonetheless you end up to solve for pi * integral of x from [0,2]

OpenStudy (anonymous):

amistre64 does the equation look like this \[ \int\limits_{2}^{0} \pi x^3/2 divided by 3/\]

OpenStudy (anonymous):

that should give you (x^2)/2 evaluated from 0 to 2 thus (2^2)/2 - (0^2)/2

OpenStudy (anonymous):

can someone just type out the work from start to finish so i undertand it bplease....:)

OpenStudy (amistre64):

the volume of a cone is 1/3 basearea times h.... because the "3" in 1/3 comes from the antiderivative of the equation for the line.... never tried it the other way around tho

OpenStudy (amistre64):

I know that the area of each slice is (S) 2pi(x^(1/2)) dx ... so ill use that:

OpenStudy (amistre64):

2pi (2/3) (x^(3/2)) is what I will use to find the volume of rotation...

OpenStudy (anonymous):

area of each slice A = pi * (Radius = sqrt x)^2

OpenStudy (anonymous):

thus A = pi*x

OpenStudy (amistre64):

4pi (2)^(2/3) 0 ----------- - --- = area under the curve 3 3

OpenStudy (amistre64):

4pi sqrt(8) --------- = area 3 fliiped my exponent around there by accident...

OpenStudy (amistre64):

8pi sqrt(2) -------- should be the area 3

OpenStudy (amistre64):

about 11.85 if I did it right, but maybe im wrong... ;)

OpenStudy (anonymous):

are you normally right...god i hope so

OpenStudy (anonymous):

for a solenoid with maximum radius 2?

OpenStudy (amistre64):

lol ..... me too :)

OpenStudy (anonymous):

ok well can you now teach me how to find the volume of the solid form by rotating the region sbout the y axis

OpenStudy (anonymous):

dude its pretty much the same

OpenStudy (amistre64):

put your equation in terms of "y" and solve it the same way

OpenStudy (anonymous):

when you rotate a graph, and want to find volume, you take the cross section of the 3D solid formed and find the equation for that section. Then you integrate that equation for cross section.

OpenStudy (anonymous):

In your given case, your y = sqrt (x) You are rotating it around x-axis

OpenStudy (anonymous):

it will give you a solenoid... kinda like a sideways bowl

OpenStudy (anonymous):

but now it need to be rotated about the y....

OpenStudy (anonymous):

i'm saying what amistre64 said is not quite right

OpenStudy (anonymous):

when you have a cross sectional area of a solenoid, it is always a circle

OpenStudy (anonymous):

and Area of that circle is A = pi * y(x)

OpenStudy (anonymous):

when you do integral of A from given bounds for x, you get the volume

OpenStudy (anonymous):

y(x)^2 sorry... i forgot the power in my A equation

OpenStudy (anonymous):

if you are rotating about Y axis, all you need to do is to put the given equation in x(y) form

OpenStudy (anonymous):

and evaluate for your range... i.e y belongs to [a, b]

OpenStudy (amistre64):

ill admit I could be wrong... so I will go "re-look" at my understanding of this :)

OpenStudy (anonymous):

ok that makes sense but what if i fine the volme of the solid ford b rotating the region about the line y= -2

OpenStudy (amistre64):

yeah, I was mistaken.... apparently I was confusing a couple a methods... for instance, take the line y=5 on the interval of [0,10]. this should produce a volume of 250pi. But if I do it the way I had in my head I get: 2pi (S) 5 dx 2pi(5x) -> 2pi5(10) = 100pi.... So I was wrong.....